OpenStudy (erik.lucken100):

differentiate e^(x/2) * sin(ax)

OpenStudy (sshayer):

\[\frac{ d }{ dx }\left( uv \right)=uv \prime+u \prime v\] \[\frac{ d }{ dx }\left( e ^{\frac{ x }{ 2 }} \right)=e ^{\frac{ x }{ 2 }}*\frac{ d }{ dx }\left( \frac{ x }{ 2 } \right)=e ^{\frac{ x }{ 2 }} *\frac{ 1 }{ 2 }\]